OpenStudy (anonymous):

find all solutions to the equation in the interval [0, 2pi). cos4x-cos2x=0 options are A) 0, 2pi/3,4pi/3 B) 0, pi/3, 2pi/3, pi, 4pi/3, 5pi/3. C) pi/6, pi/2, 5pi/6, 7pi/6, 3pi/2, 11pi/6 D) no solution

OpenStudy (turingtest):

hint:\[\cos(2x)=2\cos^2x-1\]

OpenStudy (anonymous):

cos4x=2cos^2(2x)-1; note 2x = angle thereore 2cos^2 (2x)-1-cos(2x)=0 cos(2x) [2cos(2x)-1]=0 if, cos(2x)=0 then cos^2 (x)-sin^2(x)=0 ONE OF THE SOLUTIONS is pi/4 or 45 degrees, find rest else cos(2x)= 1/2 2x= inverse cosine (1/2), 2x= pi/6, 11pi/6 therefore x= pi/12, 11 pi/12

OpenStudy (anonymous):

اHell0 cos4x=cos2x (4x=2x+2kpi ) or (4x=-2x+2k'pi) x=kpi or x= 2kpi/3 k=o x=0 k=1 x=pi k'=1 x= (2pi)/3 k'=2 x=(4pi)/3

OpenStudy (anonymous):

Thanks everybody!